# Eye-tracked Virtual Reality: A Comprehensive Survey on Methods and Privacy Challenges

### v2

@@ IV-A4 Gaze Estimation
Gaze estimation is the core focus of most eye-tracking works, as it is essential for the follow-up tasks and data processing to obtain the information on where users look. In this subsection, we do not necessarily differentiate between stationary and HMD eye tracking, as most gaze estimation algorithms are not customized for HMD use. Figure 5 broadly categorizes gaze estimation methods.
[…]
@@
According to Hansen and Ji [109], gaze estimation algorithms can be broadly classified into two categories: model-based and appearance-based. Model-based methods are sometimes also referred to as feature-based. As their name implies, model-based techniques rely on local geometric eye features such as contours and eyeball models. Model-based methods often provide high accuracy, while a common disadvantage is that they often have a limited range of operation in accordance with Zhang et al. [110], which is presumably not a central concern for HMD devices since sensors are close enough to the eyes. As surveyed by Zhang et al. [111], model-based algorithms can be further categorized into corneal reflection-based (glint-based) and shape-based (glint-free), depending on whether additional light sources are needed or not. Glint-based methods like [112, 113] often compute the location of the cornea center using the Purkinje reflection of infrared light, while a typical workflow of glint-free methods [114, 115] often begins with eye landmark detection, and then geometric features are fitted to serve gaze estimation in the downstream.
Formerly, gaze estimation methods were often heavyweight and hence could not function with eye tracking at a high rate. To address this, Feng et al. [68] introduced a novel model-based algorithm based on event-driven eye segmentation. The model tracks events (i.e., changes in pixel brightness level) to predict ROIs of near-eye images, whose resolutions reduce to 18-32% compared to original images. Then, eye region segmentation is carried out on the reduced eye images, and gaze estimation is realized upon segmentation. The model can operate at 30 Hz on a mobile device at an accuracy of 0.1-0.5. Angelopoulos et al. [69] also used motion events to promote the frame rate of gaze estimation by placing event cameras close to the eyes. While conventional model-based pupil detection algorithms approach basic pupil tracking, the motion events are utilized to update pupil location at high frequency. On an event-based eye dataset, the system reaches an accuracy of 0.45-1.75 at a frame rate of over 10 kHz. Another deficiency of most model-based algorithms is that they do not consider pupil location in 3D, which can, in turn, negatively influence eye-tracking accuracy. Lu et al. [70] proposed a solution with a 3D pupil localization model that utilizes an advanced anatomical eyeball model and accounts for the error caused by corneal reflection. With the proposed model, the authors achieved 47.5% and 18.7% error reductions for 3D pupil localization and gaze estimation, respectively, compared to prior works on their newly collected dataset, which contains ground truth of relative 3D pupil locations.
@@
An increasing shift towards appearance-based methods has recently been experienced. In contrast to model-based techniques, appearance-based methods are conceptually and structurally more straightforward: they typically use machine learning to directly learn gaze direction from photographic eye appearance (images). As surveyed by Akinyelu and Blignaut [116], prior appearance-based methods [117, 118, 119] generally relied on basic machine learning models such as linear regression, support vector machine, and random forest. Nowadays, appearance-based methods often utilize deep learning and CNNs.
@@
One example of advanced appearance-based models is by Ranjan et al. [71], which improves the robustness against head pose by head pose clustering. In contrast to former works in which one dedicated network is needed for each head pose, the authors designed a branched structure where most layers are shared while only a few final dense layers are specified for each head pose. Moreover, the authors trained the model by transfer learning from the object viewpoint estimation CNN [120] that is intrinsically more related to gaze estimation than object detection. To address data shortage, Yu and Odobez [72] suggested an unsupervised representation for gaze estimation for unlabelled eye images, and calibration works in a few-shot manner. The authors jointly trained a network for learning gaze representation and a network for gaze redirection while using warping field regularization to prevent overfitting and distortion. Unlike prior works that learn embeddings in high-dimensional space, this model captures 2D embeddings that are linked to clear physical meaning, namely, eyeball yaw and pitch. With solely 100 calibration samples, the model could reach an accuracy of 7-8.
@@
Most appearance-based gaze estimation methods only utilize static eye images and often ignore temporal traces of gaze, although they contain important information. Palmero et al. [73] demonstrated the benefit of temporal gaze sequence by implementing a many-to-one CNN-recurrent model for gaze estimation. By bringing in temporal information, the researchers realized an error reduction of up to 19.78% on average, respectively 16.91% for the horizontal axis and 23% for the vertical axis. In another promising appearance-based model, Cheng et al. [74] introduced a new asymmetric regression-evaluation network (ARE-Net) that exploits the asymmetry of human eyes. The model consists of an asymmetric regression network (AR-Net) that estimates 3D gaze direction and evaluation networks (E-Net), which assess the performance of each eye and adjust the regression strategy accordingly. While a model considering a single eye can achieve an accuracy of 6.3 on the MPIIGaze [121] dataset, the proposed method decreased the error to 5.0. Returning to the context of eye tracking in immersive virtual environments, a common problem is the HMD slippage, which can dramatically downgrade the gaze estimation accuracy of appearance-based models due to their sensitivity to camera placement. To mitigate this issue, Stojanov et al. [75] recently incorporated depth information parallel to eye appearance while using two approaches to improve the robustness against fitment and slippage. While the first approach combines features from the eye image and depth map channel-wise, the second approach exploits the so-called transformer-based cross-modal attention block [122]. This approach increases the generalizability of the model to a large extent.
@@
Besides appearance-based and model-based techniques, gaze estimation techniques combining both have also drawn attention, which are often called hybrid models. As the name implies, these methods [107, 108] commonly use deep learning models to extract eye geometric features and then map the features to gaze positions.
@@
A procedure closely related to gaze estimation is calibration, which is a prerequisite for an eye-tracking device to function correctly. However, when deployed in HMDs, calibration can be disrupted by head and body movements. Therefore, recalibration is often necessary during continuous use. There has been some work dedicated to addressing this problem. For instance, to provide a more immersive and consistent user experience, Sidenmark and Lundstroem [76] analyzed the timing and probability of fixations by exploiting the correlation between gaze location, virtual objects, and hand interaction. Based on the analyzed gaze patterns, the authors suggested an implicit and consecutive recalibration during the HMD use.
@@ IV-A5 Eye Event Classification
For data visualization and further analysis (e.g., offline analyses in psychology experiments that utilize eye tracking), eye movement is often divided into a sequence of consecutive events like fixations, saccades, and smooth pursuits. This procedure is often called eye movement event classification, eye movement segmentation, or eye movement event detection. Fixations are stable eye movements when eyes focus on a specific area or volume, whereas saccades are fast and ballistic switches between consecutive fixations. These two eye movements are often combined in the temporal dimension to create visual scanpaths. In contrast to fixations and saccades, there are some other, more fine-grained eye movements, such as smooth pursuits, defined as eye movements that follow a moving target.
In accordance with Agtzidis and Dorr [77], traditional classification methods such as [123, 124, 125], which were designed for 2D monitor-based scenarios, cannot be directly applied to eye movement event classification in VR HMDs due to the change in frame of reference. The authors suggested two approaches to alleviate such a problem: the first approach aims to transplant 2D monitor-based classification algorithms to 3D Cartesian space. At the same time, the second approach projects HMD data onto a 2D space such that traditional methods can be used without modification.
Like other gaze-related tasks, eye event classification in VR suffers from data vacancy. In a subsequent study, Agtzidis et al. [49] tried to address this issue. The authors explicitly defined primary and secondary eye events and then proposed a two-stage annotation pipeline. Furthermore, the authors developed a rule-based eye event classifier for automatic event labeling. Based on these, the authors published an annotated eye-tracking dataset for 360 videos collected from 13 observers. Recently, Rolff et al. [78] introduced a classification strategy that redefines eye event detection as a time-to-event problem. Compared to traditional algorithms that divide eye movement into discrete events, the newly proposed model can estimate the remaining time until the next saccade and predict eye behavior in the near future.
@@ IV-B Eye-based Human-Computer Interaction in VR
In the prior section, we reviewed the latest advancements in eye-tracking hardware, datasets, and algorithms that form the algorithmic fundamentals of eye-tracked VR. These advancements have opened up numerous opportunities to enhance human-computer interaction in immersive virtual environments. In this section, we comprehensively examine research in the eye-tracking-based human-computer interaction within the context of VR. We begin with exploring recent works related to eye-based interaction, explicitly addressing the challenges and potential solutions in Section IV-B1. Subsequently, we delve into predictive gaze analyses in Section IV-B2 by encompassing prediction and visualization techniques. Lastly, in Section IV-B3, we give an overview of eye-based rendering techniques that improve the real-time performance of VR devices and enhance the user experience. Figure 6 visualizes the general setup for eye tracking in VR, especially for HMDs.
@@ IV-B1 Interaction in VR
Inherent characteristics of virtual environments provide users with novel opportunities for interacting with these environments and settings. Many interaction methods require additional controllers, such as handheld ones. In contrast, gaze-based interaction can solely be executed by using built-in or integrated eye trackers in a hands-free way. However, numerous challenges exist concerning precision, time efficiency, and simplicity. In this section, we focus on works that address existing challenges around eye-based interaction techniques in VR. Table II provides an overview of all papers in this section.
TABLE II: A non-exhaustive overview of papers discussed in Section IV-B1.
@@
User interface design [126] Hand-attached menu interacted with wrist rotation and eye gaze
@@
[128] Menu design incorporating lattices as a guiding mechanism
@@
Target selection & manipulation [135] Outline Pursuits: Gaze and object movements based selection
@@
[137] EyeSeeThrough: Merging confirmation and selection processes
@@
[139] Comparison of gaze interaction in selection and drawing tasks
@@
[143] Weighted Pointer: An error-aware gaze-based interaction
[144] Fatigue analysis for eye-gaze and controller-based selection
@@
[146] Gaze+Pinch: Gaze to select targets and freehand gestures to manipulate them
@@
[151] SSVEP: Combines brain-computer interfaces and eye-tracking
[152] Keyboard visualization and dwell&click-based interaction
@@
[161] REACH+: Physical interaction in virtual spaces to enhance realism
@@
[166] Prediction for the future positions of user with LSTMs
[167] Prediction for short- and long-term future positions of user with LSTMs
@@
We first present a review focusing on the design of user interfaces based on eye tracking in VR. The main goal of these interfaces is to allow for robust interaction while minimizing the time and effort required by users. Several works have addressed this issue. For instance, Reiter et al. [126] introduced a hand-attached menu design that can be navigated with wrist rotation and interacted with the eyes, freeing up one hand for other tasks. In another work, Choi et al. [127] proposed a solution to the “Midas Touch” problem [168], a phenomenon that occurs when users’ gaze falsely interacts with an object or menu item during a search task. Choi et al. defined the gazing region between $25^{\circ}$ and $45^{\circ}$ as the “Kuiper Belt” which is not frequently targeted by the users during interaction due to being outside of the comfortable eye movement region [169, 170], but still within the physically reachable area [171, 172]. By placing menu items within the Kuiper Belt region, Choi et al. [127] showed that the number of false inputs during the search task significantly reduced, avoiding the Midas Touch problem.
@@
Kim et al. [128] suggested another menu design that aims to improve the precision of gaze-based selection and reduce the time required for selection. Incorporating lattices as a guiding mechanism for gaze gestures in the menu structure yielded fewer selection errors and shorter selection times than conventional gaze interactive menu designs [173]. Ahn et al. [129] introduced a novel gaze-based VR/AR marking menu named “StickyPie” that addressed the limitations present in existing eye-tracking interaction techniques, including overshooting and false activation. The authors reported that StickyPie, which provides a scale-invariant marking menu, achieved over 10% improvement in efficiency compared to RegularPie, a more conventional scale-variant menu design. Wei et al. [145] presented an alternative interface design for the object locating process with gaze in several VR applications with many objects, resulting in a reduced task load and time compared to traditional methods.
@@
Recently, Yi et al. [131] presented GazeDock as a view-fixed peripheral menu that is automatically displayed when the user’s gaze moves to the menu region. GazeDock, with a personalization and optimized selection algorithm, achieved an average selection time of 471 ms and a false trigger rate of 3.6% while being preferred over dwell- and pursuit-based approaches. Lee et al. [132] proposed a virtual interface, VRDoc, for a reading task that might be useful for office workers. VRDoc incorporates eye-tracking-based interaction methods to reduce document selection and navigation times and required effort compared to existing VR reading interfaces. Günther et al. [134] designed an eye-tracking-based user interface called Bionic Tracking to facilitate object tracking by exploiting smooth pursuits of eyes in a virtual environment. Based on experiments conducted with seven cell tracking experts, the interface, designed for tracking cells in 3D volumetric data using eye tracking, achieved a speed-up of 2-10 times compared to traditionally used 2D point-and-click methods.
@@
Additionally, it is essential to design user interfaces and interaction modalities by considering their specific needs and domains as well as the unique movements and conditions of the setups. To this end, Colley et al. [133] considered a vehicle setup and presented the SwiVR-Car-Seat to explore the impact of vehicle motion on VR interaction in automobiles. The authors designed an experiment using a low-cost rotating seat to observe the impact of vehicle rotation on touch, gesture, gaze, or speech interactions. The results revealed important insights for user interface design, indicating that vehicle motion had a detrimental effect on gaze- and gesture-based interaction methods. In contrast, touch and speech interactions were more resilient in such environments. Furthermore, Araujo et al. [130] proposed gaze-based control interfaces such as an overlay control interface, continuous-control interface, and waypoint navigation. In a wheelchair control setup, the semi-autonomous waypoint gaze interface yielded the fastest task completion time for each trial, implying a more favorable user experience than other interaction methods.
@@
Gaze-based user interface design requires efficient object and region selection to enable a smooth interaction experience. To this end, Sidenmark and Gellersen introduced Eye&Head [138], an approach that separately evaluates only gaze- and head-supported gaze actions for hands-free target selection. The authors found that eye-tracking data supported with head motion provides greater freedom to users during the selection process, as users’ intent is more evident in such a setup. In another work, Sidenmark et al. proposed Outline Pursuits [135], which utilizes gaze to solve object selection problems in occluded virtual scenes. The authors assigned different motions to candidate objects and analyzed the correlation between gaze and object movements. The experimental results indicated that using Outline Pursuits, the selection process requires less effort and shorter time than traditional ray-casting methods, also providing slightly better accuracy in highly occluded environments.
@@
Sidenmark et al. proposed another selection technique called Radi-Eye [136] to enable hand-free interaction in virtual smart home applications. Like the previous work, Radi-Eye uses gaze data to perform discrete or continuous target and object selections, while head movement is primarily used to confirm or modify the choice. As Radi-Eye provided a more precise and time-efficient interaction compared to previous works, it also provides important design insights for immersive virtual spaces. Qian and Teather [174] evaluated different eye-based selection methods in VR, including eye-only, eye and head combined, and head-only input methods when users were seated, and selection targets appeared in front of the participants. The authors found that the head-only selection technique was preferable during the interaction, indicating that users’ context plays an important role in the preferred interaction technique in VR.
@@
Pfeuffer et al. [146] proposed the Gaze+Pinch interaction method in VR, where the authors integrated eye tracking to select the targets and freehand gestures to manipulate them. Without relying on extra controllers, the authors’ evaluation indicated that the proposed method was intuitive for users. Liu et al. [147] focused on virtual object manipulations, particularly on rotations using gaze input, and stated that ways of interaction purely by eye gaze can be advantageous for individuals with limited use of hands and arms. Mardanbegi et al. presented EyeSeeThrough [137], a novel eye-tracking-based interaction technique that integrates both confirmation and selection processes through the line-of-sight direction. User study results demonstrated that EyeSeeThrough surpasses the performance of two-stage selection methods in terms of time and comfort. Similarly, Pfeuffer et al. [139] evaluated the effectiveness of gaze-based interaction in virtual handheld menus. The standard selection methods, including dwell time, gaze button, and cursor, were integrated with eye-tracking data and compared to a pointer-based selection method. The user studies on color selection and line drawing tasks indicated that gaze-based selection reduces physical effort compared to traditional selection methods. Minakata et al. [140] conducted a study to evaluate the performance of gaze, foot, head, and mouse pointing methods in selection tasks. The results indicated that head input was superior to gaze in terms of ease of calibration, effective target width, and throughput [175], which is the primary performance measure that considers both speed and accuracy of the users’ responses. On the other hand, gaze input performed similarly to foot input.
@@
Mutasim et al. [141] conducted another study to examine dwell and pinch movements as alternative versions of click-in devices supporting visual attention-based interaction. Their analysis indicated that pinch gestures could be a viable alternative to conventional button click-based selection methods. In another work, Mutasim et al. [142] evaluated the performance of a saccade-based selection and confirmation procedure compared to traditional approaches such as dwell and button press. The authors found that the saccade-based selection was time-efficient, yet error-prone.
@@
Recently, Sidenmark et al. [143] introduced a weighted pointer, an error-aware gaze-based interaction technique. This technique is designed to maintain stability in the presence of eye-tracker sensor errors by integrating fallback modalities. The authors demonstrated that a weighted pointer allowing automatic switching of modalities is more effective and favorable than techniques that require manual switching. In another recent work, Meng et al. [148] investigated the efficiency of hands-free text selection techniques, including selection using the dwelling, blink, and voice. The authors found that target selection using blinks outperforms the other mechanisms regarding time, accuracy, effort, and user preference.
@@
To understand practical aspects of gaze-based selection techniques, Masopust et al. [144] analyzed the fatigue occurrence, comparing selection techniques based on eye gaze and controller input. The findings revealed that prolonged use of eye gaze could lead to fatigue, making it less suitable for interactions for an extended period than controller-based techniques. Similar to Masopust et al. [144], Hou and Chen compared the eye- and controller-based selection in VR [176] and indicated that while eye-based interaction has potential, there is still some time before it becomes a mainstream interaction way in VR due to the stability and precision of the eye trackers. From another perspective, when hands-free alternatives of controller-based teleportation are evaluated, Prithul et al. [177] found that teleport activations by winking provided the most viable alternative compared to mouth gestures and dwelling, as winking led to fewer selection errors.

### v3

@@ IV-A4 Gaze Estimation
Gaze estimation is the core focus of most eye-tracking studies, as this information is essential for various follow-up tasks. In this subsection, we do not differentiate between stationary and HMD eye tracking, as most gaze estimation algorithms are not specifically tailored for HMDs. Figure 4 broadly categorizes gaze estimation methods.
Gaze Estimation Methods
@@
Gaze estimation algorithms can be broadly classified into two categories: model- and appearance-based [92]. Model-based methods are sometimes also referred to as feature-based. Model-based techniques rely on local geometric eye features such as contours and eyeball models, and they often provide high accuracy, while a common disadvantage is the limited range of operation [93], which is presumably not a central concern for HMDs since sensors are close enough to the eyes. Model-based algorithms can be further categorized into corneal reflection-based (glint-based) and shape-based (glint-free), depending on whether additional light sources are needed or not [94]. Glint-based methods often compute the location of the cornea center using the Purkinje reflection of infrared light [95], while a typical workflow of glint-free methods [96] often begins with eye landmark detection, and then geometric features are fitted to serve gaze estimation in the downstream.
Previously, gaze estimation methods were often heavyweight and therefore could not function at a high rate with eye tracking. To address this, [58] introduced a novel model-based algorithm based on event-driven eye segmentation. The model tracks events (i.e., changes in pixel brightness level) to predict ROIs of near-eye images, whose resolutions reduce to 18-32% compared to original images. Then, eye region segmentation is carried out on the reduced eye images, and gaze estimation is realized upon segmentation. The model can operate at 30 Hz on a mobile device at an accuracy of 0.1-0.5. [59] also used motion events to promote the frame rate of gaze estimation by placing event cameras close to the eyes. While conventional model-based pupil detection algorithms approach basic pupil tracking, the motion events are utilized to update pupil location at high frequency. On an event-based eye dataset, the system achieves an accuracy of 0.45-1.75 at a frame rate of over 10 kHz. Another deficiency of most model-based algorithms is that they do not consider pupil location in 3D, which can negatively influence eye-tracking accuracy. [60] proposed a solution with a 3D pupil localization model that utilizes an advanced anatomical eyeball model and accounts for the error caused by corneal reflection. The authors achieved 47.5% and 18.7% error reductions for 3D pupil localization and gaze estimation, respectively, compared to prior works.
@@
A recent shift towards appearance-based methods has been observed. In contrast to model-based techniques, appearance-based methods are conceptually and structurally simpler: they typically use machine learning to directly learn gaze direction from photographic eye appearance (images). Prior appearance-based methods [97, 98, 99] generally relied on basic machine learning, such as linear regression, support vector machine, and random forest. Nowadays, appearance-based methods often utilize deep learning and CNNs.
@@
One example of advanced appearance-based models improves the robustness against head pose by head pose clustering [61], and in contrast to former works, where a dedicated network is needed for each head pose, there is a branched structure where most layers are shared while only a few final dense layers are specified for each head pose. To address the data shortage, [62] suggested an unsupervised representation for gaze estimation for unlabelled eye images, and calibration works in a few-shot manner. The authors jointly trained two networks: one for learning gaze representation and another for gaze redirection, utilizing warping field regularization to prevent overfitting and distortion. Unlike prior works that learn embeddings in high-dimensional space, this model captures 2D embeddings that are linked to clear physical meaning, namely, eyeball yaw and pitch. With solely 100 calibration samples, the model could reach an accuracy of 7-8.
@@
Most appearance-based gaze estimation methods only utilize static eye images and often overlook temporal traces of gaze, which contain important information. [63] demonstrated the benefit of temporal gaze sequence by implementing a many-to-one CNN-recurrent model for gaze estimation. Incorporating temporal information, the researchers achieved an average error reduction of up to 19.78%, specifically 16.91% for the horizontal axis and 23% for the vertical axis. In another appearance-based model, [64] introduced a new asymmetric regression-evaluation network (ARE-Net) that exploits the asymmetry of human eyes. The model consists of an asymmetric regression network (AR-Net) that estimates 3D gaze direction and evaluation networks (E-Net), which assess the performance of each eye and adjust the regression strategy accordingly. While a model considering a single eye can achieve an accuracy of 6.3 on the MPIIGaze [100] dataset, the proposed method decreased the error to 5.0. Moving to the context of eye tracking in immersive VR, a common problem is HMD slippage, which can degrade the gaze estimation accuracy of appearance-based models due to the camera placement sensitivity. To mitigate this issue, [65] incorporated depth information parallel to eye appearance, using two approaches to improve robustness against fitting and slippage. While the first approach combines features from the eye image and depth map channels, the second approach utilizes the so-called transformer-based cross-modal attention block [101], which significantly enhances the model’s generalizability.
@@
Besides appearance-based and model-based techniques, gaze estimation techniques that combine both have also drawn attention, often referred to as hybrid models. These methods [90, 91] commonly use deep learning models to extract eye geometric features and then map the features to gaze positions.
@@
A procedure closely related to gaze estimation is calibration, which is a prerequisite for accurate eye tracking. However, when deployed in HMDs, calibration can be disrupted by head and body movements. Therefore, recalibration is often necessary during continuous use. Prior work focused on these aspects, for instance, in terms of providing a more immersive and consistent user experience, [66] analyzed the timing and probability of fixations by exploiting the correlation between gaze location, virtual objects, and hand interaction. Based on the gaze patterns, an implicit and consecutive recalibration is necessary during HMD use.
@@ IV-A5 Eye Event Classification
For data visualization and further offline analysis, eye movements are often divided into a sequence of consecutive events, like fixations, saccades, and smooth pursuits. This procedure is often referred to as eye movement event classification, segmentation, or event detection. Fixations are stable eye movements that occur when the eyes focus on a specific area or volume, whereas saccades are rapid and ballistic movements that switch between consecutive fixations. These two eye movements are combined in the temporal dimension to create visual scanpaths. In contrast to fixations and saccades, there are other, more fine-grained eye movements, such as smooth pursuits, which are defined as eye movements that follow a moving target.
Traditional classification methods [102], which were designed for 2D monitor-based scenarios, cannot be directly applied to eye movement event classification in VR HMDs due to the change in frame of reference [67]. The authors suggested two approaches to alleviate this problem: the first approach aims to transplant 2D monitor-based classification algorithms into 3D Cartesian space. At the same time, the second approach projects HMD data onto a 2D space, allowing traditional methods to be used without modification.
Like other gaze-related tasks, eye event classification in VR suffers from data vacancy. To address this, prior work explicitly defined primary and secondary eye events and then proposed a two-stage annotation pipeline [41], and developed a rule-based eye event classifier for automatic event labeling. Furthermore, [68] introduced a classification strategy that redefines event detection as a time-to-event problem, and the newly proposed strategy can estimate the remaining time until the next saccade and predict eye behavior in the near future.
@@ IV-B Eye-based Human-Computer Interaction in VR
The advancements in hardware, datasets, and algorithms have opened up numerous opportunities to enhance HCI in immersive virtual environments. In this section, we examine research in eye-tracking-based HCI within the context of VR. We begin with exploring recent works related to eye-based interaction, explicitly addressing the challenges and potential solutions in Section IV-B1. Subsequently, we delve into predictive gaze analyses in Section IV-B2 by encompassing prediction and visualization techniques. Lastly, Section IV-B3 provides an overview of eye-based rendering techniques that enhance the real-time performance of VR devices and improve the user experience. Figure 5 visualizes the general setup for eye tracking in VR, especially for HMDs.
@@ IV-B1 Interaction in VR
Inherent characteristics of virtual environments provide users with novel opportunities for interaction. Many interaction methods require additional controllers, such as handheld ones. In contrast, gaze-based interaction can be executed solely by using built-in or integrated eye trackers in a hands-free manner. However, numerous challenges exist concerning precision, time efficiency, and simplicity. In this section, we focus on works that address existing challenges around eye-based interaction techniques in VR. Table II provides an overview of papers in this section.
TABLE II: A broad overview of papers in Section IV-B1.
@@
User interface design [103] Hand-attached menu interacted with wrist and gaze
@@
[105] Menu design incorporating lattices as a guide
@@
Target selection & manipulation [112] Outline Pursuits: Gaze- and object-based selection
@@
[114] EyeSeeThrough: Merging confirmation and selection
@@
[116] Comparison of gaze in selection and drawing tasks
@@
[120] Weighted Pointer: Error-aware gaze-based interaction
[121] Fatigue analysis for gaze- and controller-based selection
@@
[123] Gaze+Pinch: Gaze for selection with gestures
@@
[128] SSVEP: Brain-computer interfaces + eye tracking
[129] Dwell&click-based interaction
@@
[138] REACH+: Physical interaction in VR to enhance realism
@@
[143] Prediction for users’ positions with LSTMs
[144] Prediction for users’ short- and long-term positions
@@
We first present a review of the design of user interfaces based on eye tracking in VR. The primary goal of these interfaces is to enable robust interaction while minimizing the time and effort required. Several works have addressed this issue. For instance, [103] introduced a hand-attached menu design that can be navigated with wrist rotation and interacted with the eyes, freeing up one hand for other tasks. [104] proposed a solution to the “Midas Touch” problem [145], a phenomenon that occurs when users’ gaze falsely interacts with an object or menu item during a search task, by defining the gazing region between $25^{\circ}$ and $45^{\circ}$ as the “Kuiper Belt” which is not frequently targeted by the users during interaction due to being outside of the comfortable eye movement region [146], but still within the physically reachable area [147, 148]. By placing menu items within the Kuiper Belt, the number of false inputs during the search task was significantly reduced, avoiding the Midas Touch.
@@
Furthermore, a different menu design that aims to improve the precision of gaze-based selection and reduce the time required for selection was suggested [105]. Incorporating lattices as a guiding mechanism for gaze gestures in the menu structure yielded fewer selection errors and shorter selection times than conventional gaze interactive menu designs [149]. [106] introduced a novel gaze-based VR/AR marking menu named “StickyPie” that addressed the limitations in existing eye-tracking interaction techniques, including overshooting and false activation. StickyPie, which provides a scale-invariant marking menu, achieved over 10% improvement in efficiency compared to RegularPie, a more conventional scale-variant menu design. [122] presented an alternative interface design for the object localization process with gaze in several VR applications featuring numerous objects, resulting in a reduced task load and time compared to traditional methods.
@@
Recently, [108] presented GazeDock as a view-fixed peripheral menu that is automatically displayed when the user’s gaze moves to the menu region. GazeDock, with a personalization and optimized selection algorithm, achieved an average selection time of 471 ms and a false trigger rate of 3.6% while being preferred over dwell- and pursuit-based approaches. [109] proposed a virtual interface, VRDoc, for a reading task that might be useful for office workers. VRDoc incorporates eye-based interaction to reduce document selection and navigation times and required effort compared to existing VR reading interfaces. [111] designed an eye-tracking-based user interface, Bionic Tracking, to facilitate object tracking by exploiting smooth pursuits in VR and achieved a speed-up of 2-10 times compared to traditionally used 2D point-and-click methods.
@@
It is also essential to design user interfaces and interaction modalities that consider the specific needs, domains, and setups. To this end, SwiVR-Car-Seat [110] explores the impact of vehicle motion on VR interaction in automobiles with an experiment using a low-cost rotating seat to observe the effect of vehicle rotation on touch, gesture, gaze, or speech interactions. The results indicate that vehicle motion has a detrimental impact on gaze and gesture interactions, whereas touch and speech interactions were more resilient. Furthermore, [107] proposed gaze-based control interfaces, including overlay control and continuous control interfaces, and waypoint navigation. In a wheelchair control setup, the semi-autonomous waypoint gaze interface yielded the fastest task completion time for each trial, indicating a more favorable user experience compared to other interaction methods.
@@
Gaze-based user interface design requires efficient object and region selection to enable a smooth interaction experience. To this end, Eye&Head [115] is an approach that separately evaluates only gaze- and head-supported gaze actions for hands-free target selection. It was found that eye-tracking data, combined with head motion, provides greater freedom to users during the selection process, as users’ intent is more evident in this setup. In another work, [112] proposed Outline Pursuits, which utilizes gaze to solve object selection problems in occluded virtual scenes. The authors assigned different motions to candidate objects and analyzed the correlation between gaze and object movements. Using Outline Pursuits, the selection process requires less effort and shorter time than traditional ray-casting methods, also providing slightly better accuracy in highly occluded environments.
@@
Another selection technique, Radi-Eye [113], enables hand-free interaction in virtual smart home applications. Like the previous work, Radi-Eye uses gaze data to perform discrete or continuous target and object selections, while head movement is primarily used to confirm or modify the choice. As Radi-Eye provided a more precise and time-efficient interaction compared to previous works, it also offers essential design insights for immersive VR. Furthermore, [150] evaluated different eye-based selection methods in VR, including eye-only, eye and head combined, and head-only input methods when users were seated, and selection targets appeared in front of the participants, and found that the head-only selection technique was preferable during the interaction, indicating that users’ context plays a vital role in interactions in VR.
@@
Gaze+Pinch [123] is another interaction method in VR, where eye movements are used for target selection and freehand gestures to manipulate them. Without relying on extra controllers, users perceived the proposed method as intuitive. [124] focused on virtual object manipulations, particularly on rotations using gaze input, and stated that ways of interaction purely by eye gaze can be advantageous for individuals with limited use of hands and arms. [114] presented EyeSeeThrough, a novel eye-based interaction technique that integrates both confirmation and selection processes through the line-of-sight direction, and EyeSeeThrough surpasses the performance of two-stage selection methods in terms of time and comfort. Similarly, [116] evaluated the effectiveness of gaze-based interaction in virtual handheld menus. The standard selection methods, including dwell time, gaze button, and cursor, were integrated with eye-tracking data and compared to a pointer-based selection method. Evaluations indicated that gaze-based selection reduces physical effort compared to traditional selection methods. Furthermore, [117] evaluated the performance of gaze, foot, head, and mouse pointing methods in selection tasks, and found that head input was superior to gaze regarding ease of calibration, effective target width, and throughput, while gaze input performed similarly to foot input.
@@
Another study examining dwell and pinch movements as alternative versions of click-in devices supporting visual attention-based interaction indicated that pinch gestures could be a viable alternative to conventional button click-based selection methods [118]. Furthermore, [119] evaluated the performance of a saccade-based selection and confirmation procedure compared to traditional approaches such as dwell and button press, and found that the saccade-based selection was time-efficient, yet error-prone.
@@
Recently, [120] introduced a weighted pointer, an error-aware gaze-based interaction technique, which is designed to maintain stability in the presence of eye-tracker errors by integrating fallback modalities. Evaluations demonstrated that a weighted pointer, allowing for automatic switching of modalities, is more effective and favorable than techniques that require manual switching. In another recent work, [125] investigated the efficiency of hands-free text selection techniques, including selection using the dwelling, blink, and voice. The target selection using blinks outperformed the other mechanisms regarding time, accuracy, effort, and preference.
@@
To understand the practical aspects of gaze-based selection, [121] analyzed the occurrence of fatigue, comparing selection techniques based on eye gaze and controller input. The findings revealed that prolonged use of eye gaze could lead to fatigue, making it unsuitable for interactions for extended periods. Similarly, [151] compared eye- and controller-based selection in VR and indicated that while eye-based interaction has potential, it still has some way to go before becoming a mainstream interaction method in VR due to the stability and precision issues with eye trackers. In another perspective, [152] evaluated hands-free alternatives of controller-based teleportation, and found that teleport activations by winking provided the most viable alternative compared to mouth gestures and dwelling, as it led to fewer selection errors.